xx CarpetEvolutionMask

CarpetEvolutionMask

Christian D. Ott, Erik Schnetter <cott@aei.mpg.de>

August 01 2005

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11enforcexmask 10Scope: privatex01 BOOLEAN

11Description: Set RHS to zero where mask value is 0?

11                    10  x01  Default: no

11

11enforcexvars  10Scope: privatex01  STRING

11Description: Enforce mask in MoL automatically on the following variables

11 Range 10 x01 Default: (none)
11                    10a list of variables or groups

11

11enlargexevolvedxregionxby 10Scope: privatex01  INT

11Description: enlarge or shrink evolved region

11 Range 10 x01 Default: (none)
110:*  10positive values enlarge the evolved region
11*:0)  10negative values shrink the evolved region. Up to bufferxwidth + nghostzones points can be removed

11

11providexbufferxmask 10Scope: privatex01  BOOLEAN

11Description: Compute mask indicating which points need only be computed for some MoL substeps

11                    10  x01  Default: no

11

11verbose  10Scope: privatex01  BOOLEAN

11Description: Do verbose debugging output

11                    10  x01  Default: no

11

11writenans  10Scope: privatex01  BOOLEAN

11Description: Write NaNs into masked regions

11                    10  x01  Default: no

11

7 Interfaces

General Implements: carpetevolutionmask

Grid Variables 7.0.1 PUBLIC GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 evolutionxmask 10 evolutionxmask x01  compact110
11                 10  x01 description11mask function
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11 10 x01 tags11prolongation=”none” InterpNumTimelevels=1
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11 bufferxmask 10 bufferxmask x01 compact110
11                 10  x01 description11buffer mask function
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11 10 x01 tags11prolongation=”none” InterpNumTimelevels=1
11                 10  x01 timelevels111
11                 10  x01variable type11INT

11

Uses header: dist.hh vect.hh carpet.hh

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/CarpetEvolutionMask. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11Conditional:
11evolutionxmask11bufferxmask

11x  11x

Scheduled Functions

CCTKxBASEGRID x evolutionmaskbasexsetupevolutionmask x set up the mask function x

11x11Type:11group

11

CCTKxPOSTREGRIDINITIAL x evolutionmaskbasexsetupevolutionmask x set up the mask function x

11x11Type:11group

11

CCTKxPOSTREGRID x evolutionmaskbasexsetupevolutionmask x set up the mask function x

11x11Type:11group

11

EvolutionMaskBasexSetupEvolutionMask x evolutionmaskbasexinitevolutionmask x initialise the mask function x

11x11Language:11c
11 x 11Options: 11global
11 x 11x  11loop-local
11 x 11Type: 11function
11 x 11Writes: 11carpetevolutionmask::evolutionxmask(everywhere)
11 x 11x  11bufferxmask(everywhere)

11

EvolutionMaskBasexSetupEvolutionMask x setupevolutionmask x set up the weight function (schedule other routines in here) x

11x11After:11evolutionmaskbasexinitevolutionmask
11 x 11Type:11group

11

SetupEvolutionMask x carpetevolutionmasksetup x set up the mask function for the restriction regions

11x11Language:11c
11 x 11Options: 11global
11 x 11x  11loop-singlemap
11 x 11Type: 11function
11 x 11Writes: 11carpetevolutionmask::evolutionxmask(everywhere)

11

MoLxPostRHS (conditional) x enforcexevolutionxmask x enforce evolution mask x

11x11Language:11c
11 x 11Reads: 11carpetevolutionmask::evolutionxmask(everywhere)
11 x 11Type: 11function